OpenStudy (anonymous):

7.) x+5y+5z=-10 x+y+z=2 x+2y+3z=-3

OpenStudy (anonymous):

how are u trying to solve this?

OpenStudy (anonymous):

substitution, linear algebra?

OpenStudy (anonymous):

well, you can do several different ways to solve this problem.

OpenStudy (anonymous):

elimination.. Please go to Vyew.com

OpenStudy (anonymous):

1. x+5y+5z = -10 2. x+y+z = 2 3. x+2y+3z = -3 I'll start with equations 1 and 2

OpenStudy (anonymous):

please go to that link to help me. its easier to understand